OpenStudy (anonymous):

Which set of side lengths cannot form a right triangle? Select one: a. 10 mm, 12 mm, 15 mm b. 9 mm, 12 mm, 15 mm c. mm, mm, mm d. 18 mm, 24 mm, 30 mm

OpenStudy (anonymous):

Hmm,what do u think? @UGG

OpenStudy (anonymous):

I don't know. :( I'm so confused.

OpenStudy (anonymous):

Hint,if there are no given value, u cannot form a right triangle

OpenStudy (anonymous):

Oh answer c didn't past c was :9/2mm, 6mm, 15/2mm

OpenStudy (anonymous):

paste*

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Thanks, and sorry about the mix up.

OpenStudy (anonymous):

use this formula \[a^2+b^2=c^2\]

OpenStudy (anonymous):

Hint,if both sides are equal,

OpenStudy (anonymous):

it is a right triangle

OpenStudy (anonymous):

let's try option A

OpenStudy (anonymous):

\[10^2+12^2=15^2\]

OpenStudy (anonymous):

\[100+144=225\]

OpenStudy (anonymous):

\[244\neq225\]

OpenStudy (anonymous):

Therefore,option A is not a right triangle

OpenStudy (anonymous):

can u do the rest? @UGG

OpenStudy (anonymous):

i will guide u

OpenStudy (anonymous):

Oh I get it now.

OpenStudy (anonymous):

Thanks so much :)

OpenStudy (anonymous):

np ^_^

OpenStudy (anonymous):

You rock :D

OpenStudy (anonymous):

u too